the area of a rectangle is 315 cm squared. the length is 6 centimeters more than the width. have to find the dimensions of the rectangle.
Answer by mananth(16949)

let width be x
length = x+6
Area of triangle = L*W
x(x+6)=315
x^2+6x=315
add-315 to both sides
x^2+6x-315=315-315
x^+6x-315=0
x^2+21x-15x-315=0
x(x+21)-15(x+21)=0
(x+21)(x-15)=0
so x = 15 OR -21
ignore negative value
width = 15
length = 21